OpenStudy (anonymous):

Compare and contrast the two quadratic equations below. In order to receive full credit, use complete sentences to describe the following: The direction each parabola opens The vertex of each parabola y = x^2 − 2x y = −2x^2 + 4x − 3

OpenStudy (anonymous):

You can tell which way the parabola opens up by looking at the sign of it's first term. The first one opens up (because the x is positive) and the second one opens down (its x is negative). To find the vertex of each parabola use this formula: vertex=-b/2a,-b^2-4ac/4a

OpenStudy (anonymous):

thank you!! lol i didnt get this at all but i kinda do now

OpenStudy (anonymous):

Think of your equation having specific parts. y=−2x^2 + 4x − 3 The general form of this equation would be: y=ax^2+bx-c Each letter stands for a term. To find the vertex of this parabola, we just use the formula: -4/2(-2),-4^2-4(-2)(-3)/4(-2) -4/-4,16-24/-8 1,-8/-8 Vertex= 1,1

OpenStudy (anonymous):

vertex=-b/2a,-b^2-4ac/4a so a =-2 b=4 c=-3 and i just fill in that equation.

OpenStudy (anonymous):

Yup. For the first parabola you don't have a C so you just fill in zero for C when you plug it into the formula.

OpenStudy (anonymous):

is the / in the formula divide ??

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\[Vertex= \frac{ -b }{ 2a },\frac{ -b^2-4ac }{ 2a }\]

OpenStudy (anonymous):

^4a

OpenStudy (anonymous):

ok now im confused is that 2 ways to find it or one big formula that u multiply to each othere??

OpenStudy (anonymous):

It's the same formula. You use it to find the vertex of ONE parabola.

OpenStudy (anonymous):

You find one first and then the other.

OpenStudy (anonymous):

for the first one i got 1

OpenStudy (anonymous):

so i got 1 in the first part and -3/2 for the second do i multiply them